# Design of an Expression Recognition Solution Employing the Global Channel-Spatial Attention Mechanism

### v2

@@ 4.4 Results
To verify the effectiveness of our method, we conducted ablation studies on each component and strategy within the method, and these studies were carried out for both single-modal data and dual-modal data. It is evident that the application of our technology significantly improved the recognition performance, and the recognition accuracy was even higher in the dual-modal scenario, which validates that the feature extraction enhancement and the unique fusion strategy can further improve the accuracy. In addition, incorporating the global channel-spatial attention mechanism can increase the accuracy by 3.3%, highlighting the importance of feature enhancement. Ultimately, through post-processing, the accuracy of the model reached 46.51%.
@@ 5 Conclusion
In this paper, we propose a method of fusing speech feature expressions to improve facial expression recognition. The first stage is called data preprocessing, and in the video processing pipeline, the preprocessing stage plays a crucial foundational role. In this stage, time masking and frequency masking techniques are first applied to perform targeted processing on the raw data, thereby filtering and optimizing the information. Subsequently, image data content is selected at equal time intervals to ensure the representativeness and regularity of the samples. The second stage is the feature extraction and enhancement stage. To address the issue of insufficient modal representation capabilities, we adopt the global channel spatial attention mechanism to enhance the features initially extracted from both the speech and image modalities, thereby improving the representation capabilities of the modalities. Finally, there is the decision-making and classification stage. A proportional decision fusion strategy is used to obtain the final emotional classification result. A coarse-to-fine granularity loss function is designed to optimize the performance of the entire network, so as to achieve more accurate dynamic facial expression recognition and classification results.

### v3

@@ 4.4 Results
According to the official competition results, the method we submitted ranked third. To verify the effectiveness of our method, we conducted ablation studies on each component and strategy within this method. These studies were carried out respectively on single-modal data and dual-modal data, and the results are shown in Table 1.It is evident that the application of our technology significantly improved the recognition performance, and the recognition accuracy was even higher in the dual-modal scenario, which validates that the feature extraction enhancement and the unique fusion strategy can further improve the accuracy. In addition, incorporating the global channel-spatial attention mechanism can increase the accuracy by 3.3%, highlighting the importance of feature enhancement. Ultimately, through post-processing, the accuracy of the model reached 46.51%.
@@ 5 Conclusion
In this paper, we propose a method of fusing audio feature expressions to improve the ability of facial expression recognition. The first stage is the data preprocessing stage. In the video processing pipeline, the preprocessing stage plays a crucial foundational role. At this stage, the frequency masking technique is first applied to conduct targeted processing on the original audio data, thus filtering and optimizing the information. Subsequently, the image data content is selected at equal time intervals to ensure the representativeness and regularity of the samples. The second stage is the feature extraction and enhancement stage. To address the issue of insufficient modal representation capabilities, we adopt the global channel-spatial attention mechanism to enhance the features initially extracted from both the audio and image modalities, thereby improving the representation capabilities of these modalities. Finally, there is the decision-making and classification stage. We adopt a decision fusion strategy based on the proportional principle to obtain the final emotional classification result. A coarse - fine granularity loss function is designed to optimize the performance of the entire network. Ultimately, we won the third place in the Expression Recognition track of the 8th Affective and Behavioral Analysis in the Wild (ABAW) Competition. In the future, we will continue to explore ways to achieve more accurate dynamic facial expression recognition and classification results.